# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v2

@@ Appendix B Dual Proximal Gradient
The condition number of the dual problem is the same as that of $\mathsf{D}\mathsf{D}^{\top}$. If we choose $\boldsymbol{\alpha}$ such that
$\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{}~{},$
we have $\|\mathsf{D}^{\top}\boldsymbol{\alpha}\|_{2}^{2}=\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
@@ Appendix C Upper Bounds on Slack Variables
$\big{\|}{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\big{\|}_{\infty}$$\leq$ $\frac{1}{|S|}\Big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\Big{)}$
$\leq$ $\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{}~{},$
@@ Appendix D Iterative Algorithm for High-Order Variational Penalties
In terms of run time, we “pay” a factor of $4^{k}$ for the number of iterations required to converge and each iteration takes $\mathcal{O}(k^{2}n)$ for solving the linear system. In comparison, when using the dual proximal gradient method, the iteration complexity grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. This implies a much inferior time complexity of $\mathcal{O}(n^{2k})$ for DPG. Pseudocode for $k>0$ is provided below.

### v4

@@ Appendix B Dual Proximal Gradient
The condition number of the dual problem is the same as that of $\mathsf{D}\mathsf{D}^{\top}$. If we choose $\boldsymbol{\alpha}$ such that $\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{},$ we have $\|\mathsf{D}^{\top}\boldsymbol{\alpha}\|_{2}^{2}=\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
@@ Appendix C Upper Bounds on Slack Variables
$\big{\|}{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\big{\|}_{\infty}$$\leq$ $\frac{1}{|S|}\Big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\Big{)}~{}\leq~{}\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{}~{},$
@@ Appendix D Iterative Algorithm for High-Order Variational Penalties
As for runtime, we “pay” a factor of $4^{k}$ for the number of iterations required for convergence where each iteration takes $\mathcal{O}(k^{2}n)$ time for solving the linear system. In comparison, when using the dual proximal gradient (DPG) method, the iteration complexity grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. This yields a substantially inferior time complexity of $\mathcal{O}(n^{2k})$ for DPG. Pseudocode for $k>0$ is provided below.